# PAC-Bayes Analysis for Recalibration in Classification

### v1

@@ Appendix C Details of experimental settings
We show the details of the datasets and the numbers of training, recalibration, and test data in Table 2. For the models, we used XGBoost [5], Random Forests [3], and a $1$-layer neural network (NN) for the KITTI, PCam, and MNIST experiments. For the CIFAR-100 experiments, we used AlexNet [16], WideResNet [37], DenseNet-BC-190 [13], and ResNeXt-29 [34]. XGBoost, Random Forests, and the $1$-layer NN were trained by adapting the code from Wenger et al. [32]  [^3] https://github.com/JonathanWenger/pycalib. Models used in the CIFAR-100 experiments were obtained from the GitHub project page: pytorch-classification (https://github.com/bearpaw/pytorch-classification).
@@
In the multiclass classification experiments, the standard method of temperature scaling [10] and the recently developed GP calibration [32] that has shown good performance were used as baselines. As a baseline method for binary classification, we additionally adopted Platt scaling [24], isotonic regression [36], Beta calibration [17], and Bayesian binning into quantiles (BBQ) [23]. We also adopted the GP calibration with mean approximation [32] in the additional experiments in Appendix D.
For our PBR method, we set the Gaussian distribution as both the posterior and prior, as in Wenger et al. [32]. We also followed the settings for training strategy, hyperparameters, and optimizers in Wenger et al. [32] to ensure a fair comparison. PBR has an optimizable parameter $1/\lambda$ in addition to the posterior parameter; thus, we selected it using grid search from the following candidates: $0\}$ , 0.01 , 0.1 , 0.25 , 0.5 , 0.75 , 0.9 , 1.0 }. The mean approximation strategy in Wenger et al. [32] can be applied for PBR. Thus, we reported the results of PBR with mean approximation (denoted as PBR app.) in Appendix D.
@@
We measured the predictive accuracy and the ECE estimated by binning, with $B=\lfloor n_{\mathrm{re}}^{1/3}\rfloor$, following our theoretical findings in Corollary 2. We conducted $10$-fold cross-validation for recalibration function training and reported the mean and standard deviation of these two performance metrics. We also used $J=100$ Monte Carlo samples from posterior $\tilde{\rho}$ to obtain $V$.
Appendix D Additional experimental results
@@
We first show all the experimental results regarding the correlation between the KL-divergence and our bound in Figures 2 and 3. We observed that the KL-divergence and generalization gap are well correlated in terms of Pearson’s and Kendall’s rank-correlation coefficients, especially in the multiclass classification tasks (MNIST and CIFAR-100). However, for the binary classification experiments, no strong correlation was observed. We conjecture that this is caused by the estimation bias of ECE through binning, which has a very slow convergence rate of $\mathcal{O}(n_{\mathrm{re}}^{1/3})$ under our optimal $B$. Since we use $n_{\mathrm{re}}=1000$, the noise due to the estimation bias can be $\mathcal{O}(1/10)$, which is larger than the ECE gap values in the binary classification experiments. Therefore, it seems that the correlation is buried in the noise.
@@
XGBoost Uncalibration $0.0119\pm 0.0004$ $0.9599\pm 0.0005$
GP calibration [32] $0.0060\pm 0.0015$ $0.9657\pm 0.0007$
GP calibration (app.) [32] $0.0067\pm 0.0017$ $0.9657\pm 0.0008$
PBR (ours) $0.0070\pm 0.0019$ $0.9657\pm 0.0006$
PBR total (ours) $\mathbf{0.0045\pm 0.0016}$ $0.9659\pm 0.0005$
PBR app. (ours) $0.0072\pm 0.0016$ $0.9655\pm 0.0006$
PBR app. total (ours) $0.0046\pm 0.0016$ $0.9659\pm 0.0005$
Platt scaling [24] $0.0076\pm 0.0018$ $0.9624\pm 0.0006$
Isotonic regression [36] $0.0079\pm 0.0023$ $0.9646\pm 0.0011$
Beta calibration [17] $0.0047\pm 0.0013$ $0.9658\pm 0.0005$
BBQ [23] $0.0071\pm 0.0025$ $0.9602\pm 0.0061$
Temperature scaling [10] $0.0090\pm 0.0018$ $0.9599\pm 0.0005$
random forest Uncalibration $0.0734\pm 0.0009$ $0.9638\pm 0.0008$
GP calibration [32] $0.0086\pm 0.0047$ $0.9628\pm 0.0015$
GP calibration (app.) [32] $0.0072\pm 0.0037$ $0.9628\pm 0.0015$
PBR (ours) $0.0090\pm 0.0048$ $0.9629\pm 0.0017$
PBR total (ours) $0.0068\pm 0.0040$ $0.9629\pm 0.0015$
PBR app. (ours) $0.0084\pm 0.0054$ $0.9629\pm 0.0017$
PBR app. total (ours) $0.0068\pm 0.0039$ $0.9629\pm 0.0016$
Platt scaling [24] $\mathbf{0.0048\pm 0.0030}$ $0.9627\pm 0.0011$
Isotonic regression [36] $0.0082\pm 0.0041$ $0.9610\pm 0.0018$
Beta calibration [17] $0.0067\pm 0.0034$ $0.9629\pm 0.0017$
BBQ [23] $0.1064\pm 0.0502$ $0.8943\pm 0.0018$
Temperature scaling [10] $0.0056\pm 0.0025$ $0.9638\pm 0.0008$
1layer NN Uncalibration $\mathbf{0.0049\pm 0.0004}$ $0.9617\pm 0.0010$
GP calibration [32] $0.0057\pm 0.0030$ $0.9682\pm 0.0006$
GP calibration (app.) [32] $0.0063\pm 0.0029$ $0.9682\pm 0.0006$
PBR (ours) $0.0074\pm 0.0037$ $0.9677\pm 0.0007$
PBR total (ours) $0.0060\pm 0.0032$ $0.9679\pm 0.0007$
PBR app. (ours) $0.0077\pm 0.0036$ $0.9678\pm 0.0007$
PBR app. total (ours) $0.0062\pm 0.0039$ $0.9680\pm 0.0007$
Platt scaling [24] $0.0089\pm 0.0039$ $0.9644\pm 0.0009$
Isotonic regression [36] $0.0081\pm 0.0037$ $0.9671\pm 0.0023$
Beta calibration [17] $0.0061\pm 0.0037$ $0.9679\pm 0.0007$
BBQ [23] $0.0316\pm 0.0732$ $0.9627\pm 0.0038$
Temperature scaling [10] $0.0092\pm 0.0042$ $0.9617\pm 0.0010$
@@
XGBoost Uncalibration $0.0516\pm 0.0009$ $0.8448\pm 0.0008$
GP calibration [32] $0.0135\pm 0.0048$ $0.8467\pm 0.0013$
GP calibration (app.) [32] $0.0140\pm 0.0047$ $0.8467\pm 0.0013$
PBR (ours) $0.0123\pm 0.0067$ $0.8467\pm 0.0012$
PBR total (ours) $0.0122\pm 0.0042$ $0.8463\pm 0.0012$
PBR app. (ours) $0.0124\pm 0.0069$ $0.8470\pm 0.0010$
PBR app. total (ours) $\mathbf{0.0121\pm 0.0049}$ $0.8464\pm 0.0013$
Platt scaling [24] $0.0407\pm 0.0103$ $0.8464\pm 0.0014$
Isotonic regression [36] $0.0167\pm 0.0052$ $0.8446\pm 0.0020$
Beta calibration [17] $0.0130\pm 0.0033$ $0.8470\pm 0.0012$
BBQ [23] $0.0178\pm 0.0079$ $0.8441\pm 0.0063$
Temperature scaling [10] $0.0542\pm 0.0070$ $0.8448\pm 0.0008$
random forest Uncalibration $0.0807\pm 0.0010$ $0.8494\pm 0.0012$
GP calibration [32] $0.0133\pm 0.0029$ $0.8485\pm 0.0017$
GP calibration (app.) [32] $0.0125\pm 0.0028$ $0.8485\pm 0.0016$
PBR (ours) $0.0121\pm 0.0044$ $0.8481\pm 0.0016$
PBR total (ours) $0.0105\pm 0.0030$ $0.8482\pm 0.0017$
PBR app. (ours) $0.0121\pm 0.0045$ $0.8481\pm 0.0018$
PBR app. total (ours) $0.0108\pm 0.0033$ $0.8481\pm 0.0020$
Platt scaling [24] $0.0109\pm 0.0034$ $0.8490\pm 0.0014$
Isotonic regression [36] $0.0159\pm 0.0049$ $0.8470\pm 0.0036$
Beta calibration [17] $0.0105\pm 0.0044$ $0.8486\pm 0.0021$
BBQ [23] $0.0491\pm 0.0092$ $0.8078\pm 0.0083$
Temperature scaling [10] $\mathbf{0.0094\pm 0.0023}$ $0.8494\pm 0.0012$
1layer NN Uncalibration $0.2063\pm 0.0012$ $0.5923\pm 0.0012$
GP calibration [32] $0.0390\pm 0.0055$ $0.6498\pm 0.0026$
GP calibration (app.) [32] $0.0394\pm 0.0048$ $0.6499\pm 0.0026$
PBR (ours) $0.0376\pm 0.0086$ $0.6509\pm 0.0015$
PBR total (ours) $0.0349\pm 0.0088$ $0.6503\pm 0.0022$
PBR app. (ours) $0.0382\pm 0.0087$ $0.6508\pm 0.0014$
PBR app. total (ours) $0.0348\pm 0.0090$ $0.6501\pm 0.0024$
Platt scaling [24] $0.0490\pm 0.0122$ $0.6220\pm 0.0059$
Isotonic regression [36] $0.0224\pm 0.0096$ $0.6476\pm 0.0041$
Beta calibration [17] $0.0274\pm 0.0072$ $0.6481\pm 0.0027$
BBQ [23] $\mathbf{0.0168\pm 0.0079}$ $0.6457\pm 0.0068$
Temperature scaling [10] $0.0256\pm 0.0178$ $0.5923\pm 0.0012$
@@
GP calibration [32] $0.0038\pm 0.0005$ $0.9785\pm 0.0007$
GP calibration (app.) [32] $0.0040\pm 0.0005$ $0.9785\pm 0.0006$
PBR (ours) $\mathbf{0.0035\pm 0.0005}$ $0.9784\pm 0.0007$
PBR total (ours) $0.0037\pm 0.0007$ $0.9785\pm 0.0007$
PBR app. (ours) $0.0038\pm 0.0002$ $0.9784\pm 0.0006$
PBR app. total (ours) $0.0041\pm 0.0007$ $0.9785\pm 0.0007$
Temperature scaling [10] $0.0055\pm 0.0017$ $0.9785\pm 0.0005$
@@
GP calibration [32] $0.0324\pm 0.0036$ $0.9663\pm 0.0011$
GP calibration (app.) [32] $0.0139\pm 0.0032$ $0.9663\pm 0.0009$
PBR (ours) $0.0052\pm 0.0018$ $0.9660\pm 0.0009$
PBR total (ours) $0.0066\pm 0.0029$ $0.9661\pm 0.0012$
PBR app. (ours) $0.0085\pm 0.0041$ $0.9662\pm 0.0009$
PBR app. total (ours) $\mathbf{0.0047\pm 0.0015}$ $0.9662\pm 0.0012$
Temperature scaling [10] $0.0062\pm 0.0009$ $0.9659\pm 0.0005$
1layer NN Uncalibration $0.0164\pm 0.0004$ $0.9760\pm 0.0005$
GP calibration [32] $0.0101\pm 0.0023$ $0.9740\pm 0.0008$
GP calibration (app.) [32] $0.0145\pm 0.0017$ $0.9740\pm 0.0009$
PBR (ours) $0.0137\pm 0.0014$ $0.9759\pm 0.0007$
@@
Temperature scaling [10] $\mathbf{0.0062\pm 0.0026}$ $0.9760\pm 0.0005$
@@
GP calibration [32] $0.2548\pm 0.0008$ $0.4373\pm 0.0009$
GP calibration (app.) [32] $0.2548\pm 0.0007$ $0.4374\pm 0.0008$
PBR (ours) $0.2548\pm 0.0007$ $0.4372\pm 0.0009$
PBR total (ours) $0.2548\pm 0.0007$ $0.4373\pm 0.0009$
@@
PBR app. total (ours) $0.2548\pm 0.0007$ $0.4374\pm 0.0008$
Temperature scaling [10] $\mathbf{0.0216\pm 0.0037}$ $0.4374\pm 0.0008$
@@
GP calibration [32] $0.0567\pm 0.0010$ $0.8129\pm 0.0011$
GP calibration (app.) [32] $0.0569\pm 0.0009$ $0.8130\pm 0.0011$
PBR (ours) $0.0504\pm 0.0054$ $0.8129\pm 0.0011$
PBR total (ours) $\mathbf{0.0349\pm 0.0093}$ $0.8129\pm 0.0011$
PBR app. (ours) $0.0569\pm 0.0007$ $0.8130\pm 0.0012$
PBR app. total (ours) $0.0569\pm 0.0008$ $0.8129\pm 0.0011$
Temperature scaling [10] $0.0374\pm 0.0028$ $0.8131\pm 0.0011$
@@
GP calibration [32] $0.0400\pm 0.0010$ $0.8229\pm 0.0014$
GP calibration (app.) [32] $0.0401\pm 0.0009$ $0.8229\pm 0.0013$
PBR (ours) $0.0320\pm 0.0050$ $0.8228\pm 0.0014$
PBR total (ours) $\mathbf{0.0310\pm 0.0082}$ $0.8230\pm 0.0014$
PBR app. (ours) $0.0399\pm 0.0024$ $0.8230\pm 0.0013$
PBR app. total (ours) $0.0402\pm 0.0008$ $0.8230\pm 0.0014$
Temperature scaling [10] $0.0401\pm 0.0019$ $0.8229\pm 0.0013$
@@
GP calibration [32] $0.0403\pm 0.0013$ $0.8230\pm 0.0014$
GP calibration (app.) [32] $0.0405\pm 0.0014$ $0.8230\pm 0.0014$
PBR (ours) $\mathbf{0.0305\pm 0.0056}$ $0.8231\pm 0.0013$
PBR total (ours) $0.0321\pm 0.0071$ $0.8230\pm 0.0014$
PBR app. (ours) $0.0384\pm 0.0032$ $0.8230\pm 0.0014$
PBR app. total (ours) $0.0369\pm 0.0033$ $0.8229\pm 0.0014$
Temperature scaling [10] $0.0420\pm 0.0025$ $0.8231\pm 0.0013$
densenet-bc-L190-k40 Uncalibration $0.0639\pm 0.0008$ $\mathbf{0.8230\pm 0.0007}$
GP calibration [32] $0.0639\pm 0.0008$ $0.8229\pm 0.0007$
GP calibration (app.) [32] $0.0640\pm 0.0008$ $0.8230\pm 0.0007$
PBR (ours) $0.0630\pm 0.0045$ $0.8228\pm 0.0006$
PBR total (ours) $0.0618\pm 0.0054$ $0.8229\pm 0.0007$
PBR app. (ours) $0.0639\pm 0.0008$ $\mathbf{0.8230\pm 0.0007}$

### v2

@@ Appendix D Details of experimental settings
We show the details of the datasets and the numbers of training, recalibration, and test data in Table 2. For the models, we used XGBoost [Chen and Guestrin, 2016], Random Forests [Breiman, 2001], and a $1$-layer neural network (NN) for the KITTI, PCam, and MNIST experiments. For the CIFAR-100 experiments, we used AlexNet [Krizhevsky et al., 2012], WideResNet [Zagoruyko and Komodakis, 2016], DenseNet-BC-190 [Huang et al., 2017], and ResNeXt-29 [Xie et al., 2017]. XGBoost, Random Forests, and the $1$-layer NN were trained by adapting the code from Wenger et al. [2020]  [^2] https://github.com/JonathanWenger/pycalib. Models used in the CIFAR-100 experiments were obtained from the GitHub project page: pytorch-classification (https://github.com/bearpaw/pytorch-classification).
@@
In the multiclass classification experiments, the standard method of temperature scaling [Guo et al., 2017] and the recently developed GP calibration [Wenger et al., 2020] that has shown good performance were used as baselines. As a baseline method for binary classification, we additionally adopted Platt scaling [Platt, 1999], isotonic regression [Zadrozny and Elkan, 2002], Beta calibration [Kull et al., 2017], and Bayesian binning into quantiles (BBQ) [Naeini et al., 2015]. We also adopted the GP calibration with mean approximation [Wenger et al., 2020] in the additional experiments in Appendix E.
Wenger et al. [2020] reported results for Platt scaling [Platt, 1999], isotonic regression [Zadrozny and Elkan, 2002], Beta calibration [Kull et al., 2017], and Bayesian binning into quantiles (BBQ) [Naeini et al., 2015], which were originally developed for binary classification tasks and adapted to the multiclass setting via a one-vs-all approach. However, for a fair comparison, our multiclass classification experiments employed temperature scaling and GP-based recalibration as baseline methods, as these approaches can be directly applied to the multiclass setting.
For our PBR method, we set the Gaussian distribution as both the posterior and prior, as in Wenger et al. [2020]. We also followed the settings for training strategy, hyperparameters, and optimizers in Wenger et al. [2020] to ensure a fair comparison. PBR has an optimizable parameter $1/\lambda$ in addition to the posterior parameter; thus, we selected it using grid search from the following candidates: $0\}$ , 0.01 , 0.1 , 0.25 , 0.5 , 0.75 , 0.9 , 1.0 }. The mean approximation strategy in Wenger et al. [2020] can be applied for PBR. Thus, we reported the results of PBR with mean approximation (denoted as PBR app.) in Appendix E.
@@
We measured the predictive accuracy and the ECE estimated by binning, with $B=\lfloor n_{\mathrm{te}}^{1/3}\rfloor$. As for the evaluation of the ECE gap, we set $B=\lfloor n_{\mathrm{re}}^{1/3}\rfloor$ following our theoretical findings in Corollary 2. We remark that we set $n_{\mathrm{te}}=n_{\mathrm{re}}$ when estimating the ECE gap. We conducted $10$-fold cross-validation for recalibration function training and reported the mean and standard deviation of these two performance metrics. We also used $J=100$ Monte Carlo samples from posterior $\tilde{\rho}$ to obtain $V$.
Appendix E Additional experimental results
@@
We first show all the experimental results regarding the correlation between the KL-divergence and our bound in Figures 3 and 4. We observed that the KL-divergence and generalization gap are well correlated in terms of Pearson’s and Kendall’s rank-correlation coefficients, especially in the multiclass classification tasks (MNIST and CIFAR-100). However, for the binary classification experiments, no strong correlation was observed. We conjecture that this is caused by the estimation bias of ECE through binning, which has a very slow convergence rate of $\mathcal{O}(n_{\mathrm{re}}^{1/3})$ under our optimal $B$. Since we use $n_{\mathrm{re}}=1000$, the noise due to the estimation bias can be $\mathcal{O}(1/10)$, which is larger than the ECE gap values in the binary classification experiments. Therefore, it seems that the correlation is buried in the noise.
@@
Finally, we evaluate the effect of using Brier and cross-entropy losses in our recalibration approach defined in Eq. (25). Specifically, we reran the experiments under the same setup described in Appendix D, and re-evaluated our two methods from Eqs. (24) and (25) in terms of ECE, classification accuracy, and cross-entropy loss.
The results are summarized in Tables 710. Our first key observation is that both of our methods—PBR and PBR total—consistently improve ECE when minimizing the Brier score. This is expected, as the Brier score upper-bounds ECE. We also observe improvements in accuracy and cross-entropy in some settings, particularly for binary classification tasks on relatively simple datasets such as KITTI and PCam.
However, when recalibration is performed using only the Brier score (with KL regularization), PBR can lead to degradation in both cross-entropy and accuracy, especially for more complex datasets and multi-class tasks. This trend is evident in Table 9 (excluding the XGBoost and Random Forest settings) and Table 10 (excluding AlexNet), and is particularly pronounced in experiments involving deep neural networks.
One possible explanation is as follows: in all experiments, the base models $f_{w}$ were trained using the cross-entropy loss. When recalibration is subsequently performed using only the Brier score, the optimization objective diverges from the training objective. This mismatch can propagate through the GP-based recalibration process, introducing inconsistencies that ultimately degrade performance in terms of accuracy and cross-entropy. Notably, this degradation is alleviated by incorporating the cross-entropy loss into the recalibration objective, as done in PBR total. This suggests that including cross-entropy helps reduce the misalignment between training and recalibration losses.
In summary, minimizing the Brier score alone does not necessarily improve cross-entropy or accuracy and may in fact worsen them. However, incorporating cross-entropy into the recalibration objective offers a more balanced trade-off, improving calibration while preserving predictive performance.
@@
XGBoost Uncalibration $0.0113\pm 0.0004$ $0.9627\pm 0.0011$
GP calibration [Wenger et al., 2020] $0.0063\pm 0.0020$ $0.9633\pm 0.0007$
GP calibration (app.) [Wenger et al., 2020] $0.0079\pm 0.0033$ $0.9640\pm 0.0013$
PBR (ours) $0.0076\pm 0.0030$ $0.9641\pm 0.0014$
PBR total (ours) $0.0063\pm 0.0014$ $0.9636\pm 0.0013$
PBR app. (ours) $0.0079\pm 0.0033$ $0.9640\pm 0.0013$
PBR app. total (ours) $0.0063\pm 0.0012$ $0.9637\pm 0.0013$
Temperature scaling [Guo et al., 2017] $0.0060\pm 0.0009$ $0.9627\pm 0.0011$
Platt scaling [Platt, 1999] $0.0076\pm 0.0014$ $0.9621\pm 0.0011$
Isotonic regression [Zadrozny and Elkan, 2002] $0.0093\pm 0.0030$ $0.9624\pm 0.0022$
Beta calibration [Kull et al., 2017] $\mathbf{0.0052\pm 0.0015}$ $0.9627\pm 0.0008$
BBQ [Naeini et al., 2015] $0.0341\pm 0.0745$ $0.9611\pm 0.0021$
Random Forest Uncalibration $0.0736\pm 0.0014$ $0.9638\pm 0.0015$
GP calibration [Wenger et al., 2020] $0.0070\pm 0.0032$ $0.9628\pm 0.0018$
GP calibration (app.) [Wenger et al., 2020] $0.0074\pm 0.0029$ $0.9626\pm 0.0018$
PBR (ours) $0.0076\pm 0.0031$ $0.9626\pm 0.0019$
PBR total (ours) $0.0068\pm 0.0033$ $0.9628\pm 0.0018$
PBR app. (ours) $0.0074\pm 0.0029$ $0.9626\pm 0.0018$
PBR app. total (ours) $0.0068\pm 0.0037$ $0.9628\pm 0.0017$
Temperature scaling [Guo et al., 2017] $0.0063\pm 0.0029$ $0.9638\pm 0.0015$
Platt scaling [Platt, 1999] $\mathbf{0.0060\pm 0.0031}$ $0.9631\pm 0.0017$
Isotonic regression [Zadrozny and Elkan, 2002] $0.0097\pm 0.0033$ $0.9612\pm 0.0024$
Beta calibration [Kull et al., 2017] $0.0063\pm 0.0023$ $0.9627\pm 0.0015$
BBQ [Naeini et al., 2015] $0.1148\pm 0.0538$ $0.8948\pm 0.0026$
1-Layer NN Uncalibration $0.0049\pm 0.0008$ $0.9619\pm 0.0014$
GP calibration [Wenger et al., 2020] $0.0047\pm 0.0008$ $0.9619\pm 0.0014$
GP calibration (app.) [Wenger et al., 2020] $0.0047\pm 0.0008$ $0.9620\pm 0.0014$
PBR (ours) $0.0047\pm 0.0008$ $0.9619\pm 0.0013$
PBR total (ours) $0.0046\pm 0.0007$ $0.9621\pm 0.0014$
PBR app. (ours) $0.0047\pm 0.0008$ $0.9619\pm 0.0014$
PBR app. total (ours) $\mathbf{0.0045\pm 0.0008}$ $0.9620\pm 0.0014$
Temperature scaling [Guo et al., 2017] $0.0104\pm 0.0050$ $0.9618\pm 0.0014$
Platt scaling [Platt, 1999] $0.0103\pm 0.0049$ $0.9642\pm 0.0016$
Isotonic regression [Zadrozny and Elkan, 2002] $0.0085\pm 0.0049$ $0.9672\pm 0.0024$
Beta calibration [Kull et al., 2017] $0.0073\pm 0.0047$ $0.9672\pm 0.0013$
BBQ [Naeini et al., 2015] $0.0324\pm 0.0760$ $0.9648\pm 0.0051$
Figure 5: $n_{\mathrm{re}}$ vs. ECE on the test dataset for experiments with the KITTI dataset. For details on our methods, we refer to Appendix D.
@@
XGBoost Uncalibration $0.0519\pm 0.0015$ $0.8445\pm 0.0015$
GP calibration [Wenger et al., 2020] $0.0122\pm 0.0053$ $0.8466\pm 0.0018$
GP calibration (app.) [Wenger et al., 2020] $0.0133\pm 0.0057$ $0.8468\pm 0.0021$
PBR (ours) $0.0139\pm 0.0043$ $0.8445\pm 0.0030$
PBR total (ours) $\mathbf{0.0113\pm 0.0043}$ $0.8464\pm 0.0024$
PBR app. (ours) $0.0140\pm 0.0042$ $0.8441\pm 0.0028$
PBR app. total (ours) $0.0116\pm 0.0055$ $0.8449\pm 0.0028$
Temperature scaling [Guo et al., 2017] $0.0605\pm 0.0048$ $0.8445\pm 0.0015$
Platt scaling [Platt, 1999] $0.0492\pm 0.0057$ $0.8474\pm 0.0011$
Isotonic regression [Zadrozny and Elkan, 2002] $0.0170\pm 0.0053$ $0.8442\pm 0.0031$
Beta calibration [Kull et al., 2017] $0.0183\pm 0.0019$ $0.8475\pm 0.0012$
BBQ [Naeini et al., 2015] $0.0143\pm 0.0098$ $0.8412\pm 0.0065$
Random Forest Uncalibration $0.0816\pm 0.0014$ $0.8495\pm 0.0018$
GP calibration [Wenger et al., 2020] $0.0124\pm 0.0027$ $0.8492\pm 0.0019$
GP calibration (app.) [Wenger et al., 2020] $0.0115\pm 0.0041$ $0.8491\pm 0.0019$
PBR (ours) $0.0109\pm 0.0040$ $0.8490\pm 0.0020$
PBR total (ours) $0.0116\pm 0.0068$ $0.8490\pm 0.0019$
PBR app. (ours) $0.0104\pm 0.0038$ $0.8491\pm 0.0020$
PBR app. total (ours) $0.0116\pm 0.0063$ $0.8489\pm 0.0019$
Temperature scaling [Guo et al., 2017] $\mathbf{0.0092\pm 0.0056}$ $0.8495\pm 0.0018$
Platt scaling [Platt, 1999] $0.0096\pm 0.0054$ $0.8496\pm 0.0014$
Isotonic regression [Zadrozny and Elkan, 2002] $0.0202\pm 0.0085$ $0.8464\pm 0.0038$
Beta calibration [Kull et al., 2017] $0.0122\pm 0.0036$ $0.8488\pm 0.0024$
BBQ [Naeini et al., 2015] $0.0529\pm 0.0079$ $0.8063\pm 0.0082$
1-Layer NN Uncalibration $0.2060\pm 0.0014$ $0.5928\pm 0.0016$
GP calibration [Wenger et al., 2020] $0.0385\pm 0.0073$ $0.6510\pm 0.0021$
GP calibration (app.) [Wenger et al., 2020] $0.0389\pm 0.0057$ $0.6507\pm 0.0023$
PBR (ours) $0.0385\pm 0.0073$ $0.6507\pm 0.0020$
PBR total (ours) $0.0384\pm 0.0060$ $0.6497\pm 0.0034$
PBR app. (ours) $0.0385\pm 0.0073$ $0.6507\pm 0.0021$
PBR app. total (ours) $0.0389\pm 0.0057$ $0.6507\pm 0.0023$
Temperature scaling [Guo et al., 2017] $0.0312\pm 0.0169$ $0.5928\pm 0.0016$
Platt scaling [Platt, 1999] $0.0484\pm 0.0105$ $0.6224\pm 0.0068$
Isotonic regression [Zadrozny and Elkan, 2002] $0.0214\pm 0.0055$ $0.6483\pm 0.0026$
Beta calibration [Kull et al., 2017] $0.0256\pm 0.0044$ $0.6478\pm 0.0026$
BBQ [Naeini et al., 2015] $\mathbf{0.0179\pm 0.0029}$ $0.6467\pm 0.0039$
Figure 6: $n_{\mathrm{re}}$ vs. ECE on the test dataset for experiments with the PCam dataset. For details on our methods, we refer to Appendix D.
@@
GP calibration [Wenger et al., 2020] $0.0038\pm 0.0006$ $0.9786\pm 0.0007$
GP calibration (app.) [Wenger et al., 2020] $0.0040\pm 0.0004$ $0.9785\pm 0.0007$
PBR (ours) $\mathbf{0.0035\pm 0.0004}$ $0.9785\pm 0.0007$
PBR total (ours) $0.0037\pm 0.0004$ $0.9785\pm 0.0006$
PBR app. (ours) $0.0036\pm 0.0003$ $0.9785\pm 0.0005$
PBR app. total (ours) $0.0036\pm 0.0004$ $0.9785\pm 0.0005$
Temperature scaling [Guo et al., 2017] $0.0055\pm 0.0017$ $0.9785\pm 0.0005$
@@
GP calibration [Wenger et al., 2020] $0.0313\pm 0.0056$ $0.9663\pm 0.0011$
GP calibration (app.) [Wenger et al., 2020] $0.0091\pm 0.0046$ $0.9641\pm 0.0016$
PBR (ours) $0.0065\pm 0.0017$ $0.9639\pm 0.0018$
PBR total (ours) $0.0057\pm 0.0028$ $0.9656\pm 0.0017$
PBR app. (ours) $0.0091\pm 0.0046$ $0.9641\pm 0.0016$
PBR app. total (ours) $\mathbf{0.0045\pm 0.0020}$ $0.9662\pm 0.0016$
Temperature scaling [Guo et al., 2017] $0.0062\pm 0.0009$ $0.9659\pm 0.0005$
1-Layer NN Uncalibration $0.0164\pm 0.0004$ $0.9760\pm 0.0005$
GP calibration [Wenger et al., 2020] $0.0101\pm 0.0023$ $0.9740\pm 0.0008$
GP calibration (app.) [Wenger et al., 2020] $0.0159\pm 0.0007$ $0.9760\pm 0.0006$
PBR (ours) $0.0137\pm 0.0014$ $0.9751\pm 0.0009$
@@
Temperature scaling [Guo et al., 2017] $\mathbf{0.0062\pm 0.0026}$ $0.9760\pm 0.0005$
Figure 7: $n_{\mathrm{re}}$ vs. ECE on the test dataset for experiments with the MNIST dataset. For details on our methods, we refer to Appendix D.
@@
GP calibration [Wenger et al., 2020] $0.2548\pm 0.0008$ $0.4373\pm 0.0009$
GP calibration (app.) [Wenger et al., 2020] $0.2548\pm 0.0007$ $0.4374\pm 0.0008$
PBR (ours) $0.2545\pm 0.0007$ $0.4374\pm 0.0008$
PBR total (ours) $0.2548\pm 0.0008$ $0.4373\pm 0.0009$
@@
PBR app. total (ours) $0.2548\pm 0.0008$ $0.4374\pm 0.0008$
Temperature scaling [Guo et al., 2017] $\mathbf{0.0216\pm 0.0037}$ $0.4374\pm 0.0008$
@@
GP calibration [Wenger et al., 2020] $0.0567\pm 0.0010$ $0.8129\pm 0.0011$
GP calibration (app.) [Wenger et al., 2020] $0.0568\pm 0.0009$ $0.8130\pm 0.0011$
PBR (ours) $0.0504\pm 0.0054$ $0.7878\pm 0.0049$
PBR total (ours) $\mathbf{0.0349\pm 0.0093}$ $0.7965\pm 0.0064$
PBR app. (ours) $0.0568\pm 0.0009$ $0.8130\pm 0.0012$
PBR app. total (ours) $0.0568\pm 0.0009$ $0.8130\pm 0.0011$
Temperature scaling [Guo et al., 2017] $0.0374\pm 0.0028$ $0.8131\pm 0.0011$
@@
GP calibration [Wenger et al., 2020] $0.0400\pm 0.0010$ $0.8229\pm 0.0014$
GP calibration (app.) [Wenger et al., 2020] $0.0400\pm 0.0009$ $0.8229\pm 0.0013$
PBR (ours) $0.0320\pm 0.0050$ $0.8023\pm 0.0084$
PBR total (ours) $\mathbf{0.0310\pm 0.0082}$ $0.8123\pm 0.0080$
PBR app. (ours) $0.0399\pm 0.0024$ $0.8167\pm 0.0047$
PBR app. total (ours) $0.0401\pm 0.0010$ $0.8229\pm 0.0014$
Temperature scaling [Guo et al., 2017] $0.0401\pm 0.0019$ $0.8229\pm 0.0013$
@@
GP calibration [Wenger et al., 2020] $0.0403\pm 0.0013$ $0.8230\pm 0.0014$
GP calibration (app.) [Wenger et al., 2020] $0.0405\pm 0.0014$ $0.8230\pm 0.0014$
PBR (ours) $\mathbf{0.0305\pm 0.0056}$ $0.8014\pm 0.0065$
PBR total (ours) $0.0321\pm 0.0071$ $0.8123\pm 0.0065$
PBR app. (ours) $0.0384\pm 0.0032$ $0.8118\pm 0.0068$
PBR app. total (ours) $0.0321\pm 0.0071$ $0.8123\pm 0.0065$
Temperature scaling [Guo et al., 2017] $0.0420\pm 0.0025$ $0.8231\pm 0.0013$
densenet-bc-L190-k40 Uncalibration $0.0639\pm 0.0008$ $0.8230\pm 0.0007$
GP calibration [Wenger et al., 2020] $0.0639\pm 0.0008$ $0.8229\pm 0.0007$
GP calibration (app.) [Wenger et al., 2020] $0.0639\pm 0.0008$ $0.8230\pm 0.0007$
PBR (ours) $0.0630\pm 0.0045$ $0.8041\pm 0.0143$
PBR total (ours) $0.0618\pm 0.0054$ $0.8195\pm 0.0078$
PBR app. (ours) $0.0639\pm 0.0008$ $0.8230\pm 0.0007$